# Dimensionality Reduction Techniques for Statistical Inference in Cosmology

### v2

@@ Abstract
We explore linear and non-linear dimensionality reduction techniques for statistical inference of parameters in cosmology. Given the importance of compressing the increasingly complex data vectors used in cosmology, we address questions that impact the constraining power achieved, such as: Are currently used methods effectively lossless? Under what conditions do nonlinear methods, typically based on neural nets, outperform linear methods? Through theoretical analysis and experiments with simulated weak lensing data vectors we compare three standard linear methods and neural network based methods. We propose two linear methods that outperform all others while using less computational resources: a variation of the MOPED algorithm we call e-MOPED and an adaptation of Canonical Correlation Analysis (CCA), which is a method uncommon in cosmology but well known in statistics. Both e-MOPED and CCA utilize simulations spanning the full parameter space, and rely on the sensitivity of the data vector to the parameters of interest. The gains we obtain are significant compared to compression methods used in the literature: up to 30% in the Figure of Merit for $\Omega_{m}$ and $S_{8}$ in a realistic Simulation Based Inference analysis that includes statistical and systematic errors. We also recommend two modifications that improve the performance of all methods: First, include components in the compressed data vector that may not target the key parameters but still enhance the constraints on due to their correlations. The gain is significant, above 20% in the Figure of Merit. Second, compress Gaussian and non-Gaussian statistics separately – we include two summary statistics of each type in our analysis.
@@ I Introduction
Yet, with the goal of extracting every bit of non-Gaussian as well as Gaussian information from cosmological observables, such as weak lensing maps, the number of widely used summary statistics (and therefore, the length of the data vectors) grows every year [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. It is crucial, then, to find suitable DR techniques so that statistical inference from the long data vectors of summary statistics can take place in a reliable and computationally efficient way. We focus here on finding DR techniques that perform best under the framework of SBI with N-body simulated data vectors. However, our results have broad applicability to cases where theory predictions are easier to obtain, or where the likelihood is Gaussian. Methodology developed here will be useful in enhancing the constraints of cosmological parameters with upcoming surveys such as the Vera C. Rubin Observatory Legacy Survey of Space and Time [34], Euclid [35], and the Nancy Grace Roman Space Telescope [36].
We consider the following linear and nonlinear DR techniques. First, we look at two linear techniques that have been widely used for decades: Principal Component Analysis (PCA; [37, 38, 39, 40, 41]) and Massively Optimized Parameter Estimation (MOPED; [42]). We also consider Canonical Correlation Analysis (CCA), which is well-known but uncommon in cosmology. Additionally, we introduce a new variation of MOPED that significantly reduces the computational cost, which we call e-MOPED. There are also three neural network based non-linear techniques that have been introduced more recently that differ in the loss functions they employ: in particular, we consider a MSE loss [43, 13, 14, 15], Variational Mutual Information Maximization [44], and Information Maximizing Neural Networks [45, 46].
@@ II.3 MOPED
Another common linear compression method is MOPED introduced in [42]. Assuming a Gaussian likelihood where the covariance matrix is independent of the parameters, the Fisher information matrix is given as
@@
Now consider the Fisher information matrix of compressed data vector $b^{T}t$ instead of $t$. Then,
@@ III.1 Neural Network with MSE
Practically speaking, since the way that the parameters affect the data vectors are non-trivially different (e.g. the effect of changes in $S_{8}$ versus changes in multiplicative biases) it is most effective to train separate neural networks for each parameter. We use Optuna [57] to optimize the network structure hyperparameters for each parameter. Let $f_{\alpha}$ be the neural network trained to predict the $\alpha^{\rm th}$ parameter. Then, the compressed data vector is
@@ III.2 Variational Mutual Information Maximization
where the sum iterates over $p_{i},\ t_{i}$.
Since we do not know the exact posterior $P({p_{i}}|{f_{\phi}(t_{i})})$, we must model the posterior with a neural density estimator with parameters $\varphi$: $\tilde{P}({p_{i}}|{f_{\phi}(t_{i})};\varphi)$. It is known that modeling the posterior with a NDE gives a lower bound of the loss function;
@@ III.3 Information Maximizing Neural Networks
Information Maximizing Neural Networks (IMNN) is another approach to DR recently introduced in [45, 46]. In essence, this is a non-linear generalization of MOPED. Instead of considering the Fisher information matrix of $b^{T}t$ as in Equation 21, we consider that of $f_{\phi}(t)$ which is a data vector $t$ compressed with a neural network $f_{\phi}$.
@@ III.4 Practicality
$\supset\sum_{i}-\frac{1}{2}(p_{i}-f_{\phi}(t_{i}))^{T}C_{\Pi}^{-1}(p_{i}-f_{\phi}(t_{i}))$
which is akin to the MSE loss with weights assigned according to the posterior covariance $C_{\Pi}$. The link between mutual information and Fisher information is often invoked to put a bound or estimate on the former using the latter in the neuroscience community. In the statistical context most relevant to our use case where the noise is Gaussian and non-negligible, one can invoke the Cramer-Rao bound to show [59, 60]:
@@
where $\operatorname{det}(F(p))$ is the determinant of the Fisher information matrix conveyed by $f_{\phi}(t)$ at parameter space location $p$. This is to say, training a neural network to increase the Fisher information matrix determinant is akin to increasing the lower bound of $I(p,f_{\phi}(t))$. IMNN indirectly accomplishes the optimization posed by VMIM, which is in turn similar to MSE optimization under realistic assumptions.
Pragmatically speaking, we were only able to obtain reliable results with VMIM when we pretrain with MSE loss and add a MSE loss term to the original VMIM loss. However, even in the best case scenario we find that the results of VMIM to be $\sim 30\%$ worse than results of NN-MSE in terms of the Figure of Merit defined in Section IV.3. Training IMNN requires the estimate of derivatives and covariances at various points in parameter space with an extremely large number of simulated data vectors which we find impractical to obtain for non-Gaussian summary statistics.
@@ IV.3 Posterior Estimation (SBI)
We estimate our posteriors using simulation-based inference (SBI). Specifically, we use neural density estimators (NDE’s) to learn the likelihood surface from our noisy mock measurements of the compressed summary statistics $c_{i}$. In particular, we employ Masked Autoregressive Flows (MAF; [79]), as implemented in the package pyDESBI [80], to estimate the conditional distribution $p(c_{i}|p_{i})$. NDE’s approximate this distribution with estimates $q(c_{i}|p_{i};\phi^{\prime})$, where the network parameters $\phi^{\prime}$ are optimized by minimizing a loss function.
@@
We evaluate each compression method for their performance using the Figure of Merit (FoM) metric, defined as
@@ V.1 Which is ‘Best’?
The performance of all other methods improves when going from WL2 to WL23_WPH_CMBWL, which is the expected behavior when adding more information to compress. However, the improvement with MOPED is very marginal. This is because the derivatives of non-Gaussian statistics computed with finite differences do not converge easily. It is possible that with even more simulations dedicated to the finite difference computation, MOPED compression of WL23_WPH_CMBWL performs better. However, as stated above, that is computationally costly.
@@ VI.1 Summary
We have studied a set of linear and nonlinear compression methods for cosmological inference problems. The linear methods are PCA, CCA, MOPED and a variation proposed here, e-MOPED (see §II). The nolinear methods are based on neural networks and are NN-MSE, VMIM and IMNN (§III). We provide theoretical arguments and experiments with simulations to identify the best methods and practices when performing DR. Our quantitative results are in the context of SBI for cosmological parameter inference though some of our findings are applicable for covariance matrix based MCMC analysis as well. We use tomographic weak lensing cosmological analysis as our application and consider a number of summary statistics: second and third moments and a set of wavelet based statistics, Wavelet Phase Harmonics and Scattering Transforms. We include the leading sources of systematic uncertainty and consider combinations of galaxy and CMB WL as well.
@@
Figure 2 summarizes the key findings in our tests with simulated data vectors. When implemented appropriately, as described in §V, none of the methods tested result in bias, so we judge them purely based on FoM. CCA and e-MOPED consistently outperform not just all other linear methods, but also NN-MSE. This is in line with other parts of machine learning literature that find linear compression to be powerful [84]. This also means we identify information loss in MOPED and NN-MSE compressions. For the task of SBI for cosmological parameter inference, CCA and e-MOPED are therefore our recommended DR methods. While PCA works well in some situations, it does poorly with low SNR data vectors, and there are a priori reasons to believe PCA and PCA-f would be unsuitable for SBI.
@@
We leave the following open questions for future investigation: Under what circumstances (or with which summary statistics) do the assumptions of e-MOPED or CCA break down enough that other methods are preferable? If the number of N-body simulations was not a limiting factor, could neural compression algorithms become the best performing? And, related, what is the precise reason for the underperformance of NN-MSE relative to the best linear methods?
@@ Appendix A PCA-f Constraints
Appendix B Correlation of $c_{i}$ and $p_{i}$
@@
One way to gauge the performance of a compression method before computing the FoM is by examining the correlation matrices of $c_{i}$ and $p_{i}$igure 9 shows |$\absolutevalue{(\rho_{c})^{\ }_{ij}}$ | and |$\absolutevalue{(\rho_{cp})^{\ }_{ij}}$ | for WL23_WPH_CMBWL using PCA, PCA-f, MOPED, e-MOPED, CCA, and NN-MSE, where $\rho_{c}$ is the auto-correlation matrix of $c_{i}$ and $\rho_{cp}$ is the cross-correlation matrix of $c_{i}$ and $p_{i}$. We show the absolute values of the elements of these correlation matrices. A good compression method should result in data vectors with uncorrelated elements so that each element adds new information, and data vectors highly correlated with the parameters of interest $S_{8}$ and $\Omega_{m}$. The first translates to a diagonal $\rho_{c}$ and the second translates to the first two columns of $\rho_{cp}$ being non-zero.

### v4

@@ Abstract
We explore linear and non-linear dimensionality reduction techniques for statistical inference of parameters in cosmology. Given the importance of compressing the increasingly complex data vectors used in cosmology, we address questions that impact the constraining power achieved, such as: Are currently used methods effectively lossless? Under what conditions do non-linear methods, typically based on neural nets, outperform linear methods? Through theoretical analysis and experiments with simulated weak lensing data vectors we compare three standard linear methods and neural network based methods. We also propose two linear methods inspired by information theory: a variation of the MOPED algorithm we call e-MOPED and an adaptation of Canonical Correlation Analysis (CCA), which is a method uncommon in cosmology but well known in statistics. Both e-MOPED and CCA utilize simulations spanning the full parameter space, and rely on the sensitivity of the data vector to the parameters of interest. In the context of our experiments, the two linear methods we propose consistently outperform the rest. The gains are significant compared to compression methods used in the literature: up to 30% in the Figure of Merit for $\Omega_{m}$ and $S_{8}$ in a realistic Simulation Based Inference analysis that includes statistical and systematic errors. We also recommend two modifications that improve the performance of all methods: First, include components in the compressed data vector that may not target the key parameters but still enhance the constraints on due to their correlations. The gain is significant, above 20% in the Figure of Merit. Second, compress Gaussian and non-Gaussian statistics separately – we include two summary statistics of each type in our analysis.
@@ I Introduction
Yet, with the goal of extracting every bit of non-Gaussian as well as Gaussian information from cosmological observables, such as weak lensing maps, the number of widely used summary statistics (and therefore, the length of the data vectors) grows every year [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. It is crucial, then, to find suitable DR techniques so that statistical inference from the long data vectors of summary statistics can take place in a reliable and computationally efficient way. We focus here on finding DR techniques that perform best under the framework of SBI with N-body simulated what the fk data vectors. However, our results have broad applicability to cases where theory predictions are easier to obtain, or where the likelihood is Gaussian. Methodology developed here will be useful in enhancing the constraints of cosmological parameters with upcoming surveys such as the Vera C. Rubin Observatory Legacy Survey of Space and Time [34], Euclid [35], and the Nancy Grace Roman Space Telescope [36].
We consider the following linear and non-linear DR techniques. First, we look at two linear techniques that have been widely used for decades: Principal Component Analysis (PCA; [37, 38, 39, 40, 41]) and Massively Optimized Parameter Estimation (MOPED; [42]). We also consider Canonical Correlation Analysis (CCA), which is well-known but uncommon in cosmology. Additionally, we introduce a new variation of MOPED that significantly reduces the computational cost, which we call e-MOPED. There are also three neural network based non-linear techniques that have been introduced more recently that differ in the loss functions they employ: in particular, we consider a MSE loss [43, 13, 14, 15], Variational Mutual Information Maximization [44], and Information Maximizing Neural Networks [45, 46].
@@ II.3 MOPED
Another common linear compression method is MOPED introduced in [42]. Assuming a Gaussian likelihood where the covariance matrix is independent of the parameters, the FIM is given as
@@
Now consider the Fisher Information Matrix (FIM) of compressed data vector $b^{T}t$ instead of $t$. Then,
@@ III.1 Neural Network with MSE
How well should we expect this to perform? Assuming that a well trained $f_{\phi}$ is an unbiased estimator of $p_{i}$ from $t_{i}$, the (co)variance of the estimate has a lower bound per the Cramer-Rao (CR) bound. Assuming Gaussian error, we have:
$f_{\phi}(t_{i})$ $=p_{i}+\epsilon_{i}$ (29)
$\epsilon_{i}$ $\sim\mathcal{N}(\vec{0},C(p_{i}))\ \textrm{and}\ C(p_{i})\geq(F(p_{i}))^{-1}$ (30)
where $C(p_{i})$ is the covariance of the estimator at $p_{i}$, $F(p_{i})$ is the FIM at $p_{i}$, and the matrix inequality $A\geq B$ implies $A-B$ is positive semidefinite. With $f_{\phi}(t)$ as the data vectors, the covariance of a posterior estimate ($C_{p}$) is bounded from below by the inverse of the FIM of the output of $f_{\phi}$. Assuming the parameter dependence of $C(p)$ is negligible, the FIM of $f_{\phi}(t)=p+\epsilon$ is equal to $C^{-1}$ since the ‘data’-parameter Jacobian is the identity; $\partial f_{\phi}(t)/\partial p=I$. In total, $C_{p}\geq C\geq F^{-1}$ in a chain of CR bounds. The philosophy behind NN-MSE is that $f_{\phi}(p)$ should perfectly capture the information content implied by the FIM.
However, there is no guarantee of this, indeed it is quite unlikely that either CR bound comes close to saturation. Furthermore, the non-linearity of $f_{\phi}$ introduces variance that is an artifact of $f_{\phi}$ and irrelevant to posterior estimation. Compressing the data by decoding the parameter value is an interpretable and human idea. However, it is not clear that parameter estimation is the best way to extract information relevant to posterior estimation. This is what motivates the approach presented in the next section - training a neural network to pick out features of the data most useful to a posterior estimator.
Practically speaking, since the way that the parameters affect the data vectors are non-trivially different it is most effective to train separate neural networks for each parameter. We use Optuna [57] to optimize the network structure hyperparameters for each parameter. Let $f_{\alpha}$ be the neural network trained to predict the $\alpha^{\rm th}$ parameter. Then, the compressed data vector is
@@ III.2 Variational Mutual Information Maximization
where the sum iterates over $p_{i},\ t_{i}$ Since we do not know the exact posterior $P({p_{i}}|{f_{\phi}(t_{i})})$, we must model the posterior with a neural density estimator with parameters $\varphi$: $\tilde{P}({p_{i}}|{f_{\phi}(t_{i})};\varphi)$. It is known that modeling the posterior with a NDE gives a lower bound of the loss function;
@@ III.3 Information Maximizing Neural Networks
Information Maximizing Neural Networks (IMNN) is another approach to DR recently introduced in [45, 46]. In essence, this is a non-linear generalization of MOPED. Instead of considering the FIM of $b^{T}t$ as in Equation 21, we consider that of $f_{\phi}(t)$ which is a data vector $t$ compressed with a neural network $f_{\phi}$.
@@ III.4 Practicality
$\supset\sum_{i}-\frac{1}{2}(p_{i}-f_{\phi}(t_{i}))^{T}C_{p}^{-1}(p_{i}-f_{\phi}(t_{i}))$
which is akin to the MSE loss with weights assigned according to the posterior covariance $C_{p}$. The link between mutual information and Fisher information is often invoked to put a bound or estimate on the former using the latter in the neuroscience community. In the statistical context most relevant to our use case where the noise is Gaussian and non-negligible, one can invoke the Cramer-Rao bound to show [59, 60]:
@@
where $\operatorname{det}(F(p))$ is the determinant of the FIM conveyed by $f_{\phi}(t)$ at parameter space location $p$. This is to say, training a neural network to increase the FIM determinant is akin to increasing the lower bound of $I(p,f_{\phi}(t))$. IMNN indirectly accomplishes the optimization posed by VMIM, which is in turn similar to MSE optimization under realistic assumptions.
Pragmatically speaking, we were only able to obtain reliable results with VMIM when we pretrain with MSE loss and add a MSE loss term to the original VMIM loss. However, even in the best case scenario we find that the results of VMIM to be $\sim 30\%$ worse than results of NN-MSE in terms of the Figure of Merit (FoM) defined in Section IV.3. Training IMNN requires the estimate of derivatives and covariances at various points in parameter space with an extremely large number of simulated data vectors which we find impractical to obtain for non-Gaussian summary statistics.
@@ IV.3 Posterior Estimation (SBI)
We estimate our posteriors using simulation-based inference (SBI). Specifically, we use neural density estimators (NDE’s) to learn the likelihood surface from our noisy mock measurements of the compressed summary statistics $c_{i}$. In particular, we employ Masked Autoregressive Flows (MAF; [79]), as implemented in the package pyDELFI [80], to estimate the conditional distribution $p(c_{i}|p_{i})$. NDE’s approximate this distribution with estimates $q(c_{i}|p_{i};\phi^{\prime})$, where the network parameters $\phi^{\prime}$ are optimized by minimizing a loss function.
@@
We evaluate each compression method for their performance using the FoM metric, defined as
@@ V.1 Which is ‘Best’?
The performance of all other methods improves when going from WL2 to WL23_WPH_CMBWL, which is the expected behavior when adding more information to compress. However, the improvement with MOPED is very marginal so MOPED falls drastically behind other linear methods. This is because the derivatives of non-Gaussian statistics computed with finite differences do not converge easily. It is possible that with even more simulations dedicated to the finite difference computation, MOPED compression of WL23_WPH_CMBWL performs better. However, as stated above, that is computationally costly.
A toy exercise indicates that CCA and e-MOPED are more effective than NN-MSE when the response to parameter is linear within the simulation suite prior. The details of the toy exercise can be found in Appendix B. The results here indicate that WL23_WPH_CMBWL considered here falls within that regime of linearity.
@@ VI.1 Summary
We have studied a set of linear and non-linear compression methods for cosmological inference problems. The linear methods are PCA, CCA, MOPED and a variation proposed here, e-MOPED (see §II). The nolinear methods are based on neural networks and are NN-MSE, VMIM and IMNN (§III). We provide theoretical arguments and experiments with simulations to identify the best methods and practices when performing DR. Our quantitative results are in the context of SBI for cosmological parameter inference though some of our findings are applicable for covariance matrix based MCMC analysis as well. We use tomographic weak lensing cosmological analysis as our application and consider a number of summary statistics: second and third moments and a set of wavelet based statistics, Wavelet Phase Harmonics and Scattering Transforms. We include the leading sources of systematic uncertainty and consider combinations of galaxy and CMB WL as well.
@@
Figure 2 summarizes the key findings in our tests with simulated data vectors. When implemented appropriately, as described in §V, none of the methods tested result in bias, so we judge them purely based on FoM. CCA and e-MOPED consistently outperform not just all other linear methods, but also NN-MSE given our setup. This also means we identify information loss in MOPED and NN-MSE compressions. For the task of SBI for cosmological parameter inference, CCA and e-MOPED are therefore our recommended DR methods. While PCA works well in some situations, it does poorly with low SNR data vectors, and there are a priori reasons to believe PCA and PCA-f would be unsuitable for SBI. MOPED is severely limited by the difficulty of numerical derivative convergence. Toy exercise explored in Appendix B reveal that the relevant problem is in a regime where NN-MSE is inadequate.
At the same time, given the non-linear relationship between our parameters and data vectors, it is very plausible that there is a neural network architecture and loss function that out-perform CCA and e-MOPED. We find VMIM and IMNN to be impractical, and NN-MSE to underperform, but there may be a neural DR technique that is practical and powerful. This is especially the case for levels of noise and non-linearity beyond those considered here.
@@
We leave the following open questions for future investigation: Under what circumstances (or with which summary statistics) do the assumptions of e-MOPED or CCA break down enough that other methods are preferable? Is there a neural compression algorithm that would out-perform e-MOPED or CCA? How do the conclusions here translate to future surveys?
@@ Appendix A PCA-f Constraints
Appendix B Toy Model
Figure 9: Results from a toy model of posterior estimation in varying linearity and noise level regimes using CCA, e-MOPED, and NN-MSE. The figures reflect averages over 20 realizations of NDEs. When noise levels are high, the 3 methods perform equivalently well. Otherwise, linear methods are better than NN-MSE when the response to the parameters is linear.
We define a toy model of a posterior estimation problem to identify the regimes in which CCA and e-MOPED might out-perform NN-MSE as we find with weak lensing data vectors.
First, we define a 6 dimensional parameter space. The first two parameters will be the parameters of interest, and the rest nuisances. We uniformly sample the $(0,1)^{6}$ hypercube to populate our simulation prior with $N_{\rm sim}=20000$ points. Half of the points were used for the compression step, the other half was used for the SBI after compression. This mimics realistic $N_{\rm sim}$ and parameter space for weak lensing SBI scenarios.
Then we define 100 dimensional linear and non-linear data vectors. For the linear data vector, we define an arbitrary ${100\times 6}$ matrix $A$ where every element is between $-1$ and $1$ such that $Ap$ is the noiseless linear data vector where $p$ is a given parameter vector. For the non-linear data vector, we define an arbitrary ${100\times 6}$ matrix $B$ where every element is between $2$ and $6$. Then, the $i^{\rm th}$ element of the noiseless non-linear data vector $v$ is defined as
$v_{i}=\sum_{j=1}^{6}p_{j}^{B_{i,j}^{\ }}$ (44)
where $p_{j}$ is the $j^{\rm th}$ parameter.
The noise in our analysis was modeled as additive white noise, meaning it was independent across data vector components and parameters, with a defined standard deviation. For linear data vectors, we used two noise levels: low noise with $\sigma_{\rm low}=0.02$ and high noise with $\sigma_{\rm hi}=0.2$. When working with non-linear data vectors, we chose their $\sigma_{\rm low}$ and $\sigma_{\rm hi}$ values to ensure that the determinant of the Fisher Information Matrix (calculated for the first two parameters) matched the corresponding values from the linear case at each noise level. This approach allowed us to maintain consistent noise conditions when comparing linear and non-linear models. Practically speaking, this is just a constant rescaling of $\sigma_{\rm low},\sigma_{\rm hi}$ as a function of $A$ and $B$. Note that this makes the noise levels fair at the level of the Cramer-Rao bounds, and does not guarantee comparable FoM.
With these data vectors, we use CCA, e-MOEPD, and NN-MSE to perform SBI on the first two parameters given a data vector generated at $p_{\rm fid}=(0.5,0.5,0.5,0.5,0.5,0.5)$. The resulting FoM are shown in Figure 9. We find that when noise levels are high, the choice of DR is not very important. We also find that when noise levels are low, the choice of DR impacts FoM. If the response to parameter is non-linear, NN-MSE is clearly better. On the other hand, when the response to parameter is linear, linear methods out-perform NN-MSE.
An important note here is that response to parameter is always linear given a small enough simulation suite prior. Even if the true response is non-linear (as is the case with $\sigma_{8}$ and WPH, small scale WL2, etc.), if the simulation suite prior is small enough the response to parameter is linear to dominant order.
This gives some context to our findings that CCA and e-MOPED out-perform NN-MSE. This toy exercise would suggest that even with the non-Gaussian statistics in WL23_WPH_CMBWL, DES Y3 level weak lensing SBI is in a regime where
  1. 1.
    noise levels are low enough for choice of DR to matter, and
  2. 2.
    the response to parameter is linear enough within the simulation suite prior that linear methods of DR are better than NN-MSE.
From this one might speculate that with future surveys with lower noise, the choice of DR will become even more important. Furthermore, the Bayesian nature of cosmology affords future surveys tighter simulation suite priors, making linear methods even more of valid choice over NN-MSE.
Appendix C Sensitivity to Scale Cuts
Figure 10: FOM obtained with WL23_WPH_CMBWL while varying the number of scales cut from the smallest scales. The curves reflect the mean of 20 NDE realizations, and the error bars are the standard error. The conclusion that CCA (blue) and e-MOPED (orange) are the best out of the methods showcased is robust to scale cuts. NN-MSE (green) and MOPED (red) fail to efficiently capture the information at the smallest scale due to the low SNR.
To probe the sensitivity of our results to scale cuts, we repeated the exercise of computing the FOM in $\Omega_{m}-S_{8}$ plane with WL23_WPH_CMBWL using CCA, e-MOPED, NN-MSE, MOPED, and PCA while varying the number of smallest scales cut. We stop at cutting the two smallest scales, because further cuts are too restrictive for the WPH portion of the data vector.
We find that CCA and e-MOPED (in that order) outperform all other methods across different scale cuts. The bend in the NN-MSE curve reflects the fact that the smallest scale is too noisy for neural networks to efficiently extract information. FOM from MOPED appears to increase with the first scale cut, because, as mentioned elsewhere, the performance of MOPED hinges on the convergence of numerical derivatives. Numerical derivatives of non-Gaussian statistics at the noisiest scales is extremely noisy and renders MOPED ineffective overall. This highlights the point that NN-MSE and MOPED are ill suited to compressing low SNR statistics.
Appendix D Correlation of $c_{i}$ and $p_{i}$
@@
One way to gauge the performance of a compression method before computing the FoM is by examining the correlation matrices of $c_{i}$ and $p_{i}$igure 11 shows |$\absolutevalue{(\rho_{c})^{\ }_{ij}}$ | and |$\absolutevalue{(\rho_{cp})^{\ }_{ij}}$ | for WL23_WPH_CMBWL using PCA, PCA-f, MOPED, e-MOPED, CCA, and NN-MSE, where $\rho_{c}$ is the auto-correlation matrix of $c_{i}$ and $\rho_{cp}$ is the cross-correlation matrix of $c_{i}$ and $p_{i}$. We show the absolute values of the elements of these correlation matrices. A good compression method should result in data vectors with uncorrelated elements so that each element adds new information, and data vectors highly correlated with the parameters of interest $S_{8}$ and $\Omega_{m}$. The first translates to a diagonal $\rho_{c}$ and the second translates to the first two columns of $\rho_{cp}$ being non-zero.